Two free positive charges 4q and q are a distance l apart. What charge Q is needed to achieve
equilibrium for the entire system and where should it be placed form charge q? Why we equate equilibrium at Q rather than at q or 4q

Grade:12

1 Answers

guddu gupta
3 years ago
You can apply equilibrium on any of the three charges, but usually, for simplicity, we use the Q. The Q must be close to the smaller charge.
Let the distance from q be x
k(q)(Q)/x2= k(4q)(Q)(l-x)2
x=l/3